2. (10 points) Find the knot vector for a quartic (k = 5) B-spline created from ten control points.
Answer #1

t (titk-0) Nitlok+). Given that 3, KLU) ki; Ni,k(0: (U-ti) Ninkai Since generat expression for Basic function is k=5 0+1=10 =Now K= 5, n=9 tilo zis 14) to o OL5 from ti=0 125 @simi looly tzétz= ty: 0 5566 9 t5-5-5+1 = 1 from t6=6-5+1=2 ... 55649 tg=t